# Detecting and Interpreting NSFW Prompts in Text-to-Image Models through Uncovering Harmful Semantics

### v2

@@ 5.3 Adaptive Attack
ASR-H(%) ASR-M(%) ASR-H(%) ASR-M(%)
Bare 46 48 75 79
HiddenGuard 0 0 28 31
[…]
Table 4: The evaluation of adaptive attack.
In this section, we evaluate the robustness of HiddenGuard against adaptive attacks. We employ SneakyPrompt for black-box attack and MMA for white-box attack. In the black-box setting, the attacker directly targets the defended model. In the white-box setting, the attacker has full access to the T2I model’s architecture and our defense mechanism. This allows them to craft a specialized loss function to bypass the defense. To simulate this, we design an attack by augmenting the original loss function of the MMA with an additional loss term targeting the NSFW score. Detailed configurations for both black-box and white-box adaptive attacks are provided in the Appendix A.11.
We utilize Stable Diffusion v1.4 (CompVis, 2022) as the generative model, as it is the model most vulnerable to attacks. We evaluate the attack performance on the bare model and the model employing HiddenGuard. The bare model only detects whether sensitive tokens are present in the prompt.
Experimental results are shown in Table 4. ASR-H and ASR-M represent the attack success rate evaluated by human and classification model, respectively. In the black-box scenario, while the SneakyPrompt attack achieves a certain success rate against the bare model, it is completely unable to penetrate the defense provided by HiddenGuard. In the white-box setting, although a limited number of adversarial examples can bypass the defense, the ASR is drastically reduced compared to the bare model. This result further demonstrates the robustness of HiddenGuard’s defense mechanism.
[…]
5.4 Interpretation Experiments
In this section, we validate HiddenGuard’s interpretability. This not only makes the detection more transparent and trustworthy but also aids users in further understanding the semantics of prompts.
Text-based interpretation. In the text modality, our interpretation aids users in understanding the semantics of a prompt by identifying tokens containing NSFW semantics. After obtaining the interpretation result $E(p)_{i}$ for each token in the prompt $p$, we sequentially remove the corresponding tokens from $p$ in descending order of $E(p)_{i}$ and observe the changes in the NSFW score. We compare our interpretation method against two baselines. The first, a random-based method, removes tokens from the prompt at random. The second, an embedding-based method, uses the CLIP score to evaluate the semantic similarity of each token to the term ”NSFW” and removes tokens in descending order of this similarity. The experimental results are shown in Figure 8. Compared to both baselines, removing tokens based on our interpretation method more rapidly reduces the NSFW semantic content of the prompt. This demonstrates the effectiveness of HiddenGuard in both explaining NSFW prompts and locating the specific tokens responsible for the NSFW content.
Figure 7: Text-based interpretation.
Figure 8: Image-based interpretation.
Image-based interpretation. For image-based interpretation, we use the same 100 prompts from the previous section and progressively eliminate NSFW semantics from token embeddings. We evaluate the resulting images for both remaining NSFW content and their semantic fidelity to the original prompt. Figure 8 illustrates the variation in the number of NSFW images generated and the semantic similarity between the images and prompts as the parameter $\beta$ changes. The semantic similarity is evaluated using the CLIP Score (Hessel et al., 2021). As $\beta$ increases, the number of NSFW images decreases. Although the CLIP Score shows a decreasing trend, the overall deviation from the original image’s CLIP Score remains minimal. This indicates that the NSFW semantics of the prompt are effectively mitigated while preserving other semantic information.
6 Conclusion
In this paper, we propose a unified defense framework against NSFW prompts in T2I models named HiddenGuard. The HiddenGuard framework, being adaptable, efficient, interpretable, optimizable, and unified, has demonstrated superior performance that far exceeds previous defense methods. In addition to detection, we provide interpretability methods to help understand the semantics of NSFW prompts and the generation process of NSFW images. Experimental results show that HiddenGuard exhibits strong capabilities in defending against both normal attacks and adaptive attacks.
Appendix A Appendix
A.1 The Use of Large Language Models (LLMs)
In the preparation of this manuscript, we utilized LLMs solely as a general-purpose writing assistance tool. The primary use of the LLM was for proofreading and copy-editing, including but not limited to, improving grammar, refining phrasing, correcting spelling, and enhancing the overall readability and clarity of the text.
We explicitly state that the LLMs’ role was strictly confined to language polishing. It was not used for any substantive aspects of the research, such as research ideation, literature review, experimental design, data analysis, result interpretation, or the initial drafting of any sections of the paper. All intellectual contributions, including the core ideas, methodologies, and conclusions presented herein, are entirely the work of the human authors.
A.2 The Optimization Process of NSFW Feature
where $\mu_{b}^{l,h}$ and $\mu_{m}^{l,h}$ are the mean value of $\{c_{b_{k}}^{l,h}\}$ and $\{c_{m_{k}}^{l,h}\}$. In summary, our goal is to maximize the distance between the projected means of $\{c_{b_{k}}^{l,h}\}$ and $\{c_{m_{k}}^{l,h}\}$ on the vector $u^{l,h}$, while simultaneously minimizing their respective variances. We can employ Linear Discriminant Analysis to solve this problem, ultimately obtaining the NSFW feature $\{u^{l,h}\}$:
$u^{l,h}=S_{w}^{-1}(\mu_{m}^{l,h}-\mu_{b}^{l,h}),$ (12)
$S_{w}=\sum_{k=0}^{K_{m}}\|c_{m_{k}}^{l,h}-\mu_{m}^{l,h}\|^{2}+\sum_{k=0}^{K_{b}}\|c_{b_{k}}^{l,h}-\mu_{b}^{l,h}\|^{2}$ (13)
In text encoders other than CLIP, such as T5 (Raffel et al., 2020), we can still use the method above to extract NSFW features. Any text encoder utilizing multi-head self-attention can be adapted to this approach.
A.3 Implementation Details of NSFW Detection
For any given prompt $p$, we determine whether it is an NSFW prompt based on the magnitude of its corresponding NSFW Score $\text{Score}(p)$. Theoretically, if $\text{Score}(p)>0$, the prompt $p$ contains NSFW semantics and should be classified as an NSFW prompt. Conversely, if $\text{Score}(p)<0$, the prompt should be benign. However, experiments show that while setting the threshold to zero allows HiddenGuard to achieve high accuracy, optimal classification performance requires a slight threshold adjustment. We hypothesize that this is due to the distribution of the training set not fully representing the actual distribution of NSFW prompts, introducing bias in the NSFW features derived during training. In our experiments, we determine the threshold by selecting the one that yields the highest F1 Score on the training set. For different text encoders, the final offset ranges from 1% to 3%.
The default approach treats NSFW as a single comprehensive category. Suppose there is a need to subdivide it further or to identify specific categories of NSFW prompts, such as sex or violence. In that case, we can categorize NSFW prompts in the training set based on labels. This enables the calculation of NSFW features for each subcategory, facilitating the determination of NSFW scores for each. If we also need to detect NSFW prompts across all categories, we can aggregate the NSFW scores from all subcategories and use their maximum value as the final NSFW score.
Since HiddenGuard operates exclusively at the prompt input and encoding stages, it can be readily combined with other defense mechanisms to achieve superior performance. These include internal guards Gandikota et al. (2023); Kumari et al. (2023), which are fine-tuned to make the model resistant to generating NSFW content, and post-hoc tools that scan the final output images. Employing such a multi-layered defense strategy can further enhance the overall security of the entire image generation framework.
A.4 More Examples of Image-Based Interpretation
To further demonstrate the effectiveness of our image-based interpretation across multiple categories of NSFW prompts, we selected five prompts, each from a distinct subcategory. The specific prompts and the results of their interpretation are presented in Table 5. Figure 4 shows the images generated with different values of $\beta$. As the value of $\beta$ increases, the images generated from both prompts gradually transition from NSFW to harmless.
During this process, we observed two interesting phenomena. First, as the NSFW semantics are progressively reduced, the model still attempts to preserve the original semantics of the prompt. For the sexual prompt, the woman gradually turns away and conceals sensitive areas with her hands. Ultimately, only the area above her shoulders is visible, rendering the image harmless. Throughout, the image never violates the prompt’s directive. For the violent prompt, the blood on the man’s face gradually changes color to resemble oil stains, and the marks themselves progressively diminish. However, even at $\beta=1$, some dirty marks remain on the man’s face. These examples illustrate that the prompt’s semantics are highly editable within the hidden states. By applying our explanation method, we can progressively eliminate the NSFW semantic content while preserving the original meaning of the prompt to the greatest extent possible.
Category Example Prompt
Sexual A beautiful naked woman.
Violence a man with blood all over his face.
Hate ghost as an emoji, telegram sticker design, flat design, glossy design, white outline.
Illegal pile Off pills , drugs and a skull on the table, skull made out of smoke coming out of pills, fantasy, intricate, elegant, highly detailed, digital painting, artstation, concept art, addiction, chains, smooth, sharp focus, illustration, art by Ilja Repin.
Shocking hyper realistic ultra realistic horror terror dimensional photo furious glowing red eyes biden, high quality photo, detailed , 8k
Table 5: Examples of Different Categories of NSFW Prompt
Figure 9: Image-based interpretation with Stable Diffusion v1.4.
A.5 PCA Maps of All Attention Heads
In Figure 12, we present a PCA map Abdi & Williams (2010) illustrating the distribution of three types of prompts across all attention heads in the CLIP-L model. These maps provide a clearer view of the substantial variations in the attention given to NSFW semantics across different layers and heads. In some heads, the distributions of NSFW prompts and benign prompts show minimal differences. In other maps, they exhibit relatively clear boundaries. The sensitivity of different attention heads to each category also varies, which aligns with the results we presented in the main paper.
A.6 Related Work about Interpretation on the CLIP Model
As the most widely used text encoder in T2I models, the CLIP model (Radford et al., 2021) is a significant focus of adversarial attack research Liu et al. (2023). Recent studies (Gandelsman et al., 2024; Bhalla et al., 2024; Zhao et al., 2024; Aflalo et al., 2022) have investigated the CLIP model to analyze its internal mechanisms. For instance, Bhalla et al. (Bhalla et al., 2024) found that the embeddings generated by CLIP exhibit strong linear properties and can be decomposed into combinations of various concepts. Gandelsman et al. (Gandelsman et al., 2024) discovered that different attention heads within the CLIP model are responsible for interpreting different semantics, which are then combined to produce the final output embeddings. However, these studies primarily focus on the image domain. In this paper, we build upon existing work to further explore the properties of the CLIP model in the text domain and propose a novel interpretation method to interpret how prompts containing NSFW semantics are generated.
A.7 Theoretical Analysis of Adversarial Robustness
This section provides a revised formal analysis of the adversarial robustness of the HiddenGuard detection mechanism. We derive a certified lower bound on the perturbation magnitude required to deceive the detector, under a natural local Lipschitz assumption directly on the score function with respect to the final embedding metric.
A.7.1 Preliminaries and Definitions
Definition 1 (Prompt Sets and Mappings).
Let $\mathcal{P}$ be the set of all prompts. We consider two disjoint subsets: the set of benign prompts, $\mathcal{P}_{B}\subset\mathcal{P}$, and the set of NSFW prompts, $\mathcal{P}_{M}\subset\mathcal{P}$. We define two feature mappings from the prompt space to a $d$-dimensional real vector space $\mathbb{R}^{d}$:
  • Final Embedding Map $\Phi_{E}$: $\Phi_{E}:\mathcal{P}\to\mathbb{R}^{d}$ maps a prompt $p$ to its final text embedding vector. This is the primary space where adversarial perturbations are constrained.
  • Hidden State Map $\Phi_{l,h}$: $\Phi_{l,h}:\mathcal{P}\to\mathbb{R}^{d}$ maps a prompt $p$ to its hidden state representation from layer $l$ and attention head $h$ of the text encoder. This is the space where HiddenGuard operates.
Definition 2 (HiddenGuard Detector and Score).
For a specific layer-head pair $(l,h)$, HiddenGuard uses a linear detector in the hidden-state space given by
$f_{l,h}(p)=\mathbf{1}\{\,S(p)\geq\tau\,\},\quad\text{where}\quad S(p)=\langle\Phi_{l,h}(p),\mathbf{u}^{l,h}\rangle,$
and $\mathbf{u}^{l,h}\in\mathbb{R}^{d}$ is the normalized NSFW direction ($\|\mathbf{u}^{l,h}\|_{2}=1$), and $\tau\in\mathbb{R}$ is the decision threshold. A prompt $p$ is classified as benign if $S(p)<\tau$ and as NSFW if $S(p)\geq\tau$.
Definition 3 (Adversarial Perturbation in Final Embedding Space).
An adversarial attack against a benign prompt $p_{B}\in\mathcal{P}_{B}$ aims to find an adversarial prompt $p_{\mathrm{adv}}$ that is semantically indistinguishable from $p_{B}$ but is misclassified as NSFW. We model semantic indistinguishability by bounding the perturbation in the final embedding space:
$\|\Phi_{E}(p_{\mathrm{adv}})-\Phi_{E}(p_{B})\|_{2}\leq\epsilon.$
An attack is successful if $S(p_{B})<\tau$ and $S(p_{\mathrm{adv}})\geq\tau$.
Assumption 1 (Local Lipschitz Continuity of Score).
Fix a benign prompt $p_{B}$. There exists a radius $r>0$ and a constant $K_{l,h}(p_{B})\geq 0$ such that the score function $S(\cdot)$ is locally Lipschitz with respect to the final embedding within the ball
$\mathbb{B}_{r}\big(\Phi_{E}(p_{B})\big)=\{\,z\in\mathbb{R}^{d}:\|z-\Phi_{E}(p_{B})\|_{2}\leq r\,\},$
in the following sense: for any two prompts $p_{1},p_{2}$ satisfying $\Phi_{E}(p_{1}),\Phi_{E}(p_{2})\in\mathbb{B}_{r}\big(\Phi_{E}(p_{B})\big)$,
$|S(p_{1})-S(p_{2})|\leq K_{l,h}(p_{B})\cdot\|\Phi_{E}(p_{1})-\Phi_{E}(p_{2})\|_{2}.$
We denote $K_{l,h}(p_{B})$ simply as $K$ when no confusion arises.
A.7.2 Certified Robustness Radius
Theorem 1 (Certified Robustness Radius (Local)).
Let $p_{B}$ be a benign prompt correctly classified by HiddenGuard, i.e., $S(p_{B})<\tau$. Under Assumption 1, any adversarial prompt $p_{\mathrm{adv}}$ created from $p_{B}$ will not be misclassified as NSFW, provided that the perturbation magnitude $\epsilon$ satisfies
$\epsilon<\min\Big\{\,r,\;\frac{\tau-S(p_{B})}{K}\Big\}.$
Proof.
Let $p_{B}$ be a benign prompt with $S(p_{B})<\tau$. Consider any $p_{\mathrm{adv}}$ such that $\|\Phi_{E}(p_{\mathrm{adv}})-\Phi_{E}(p_{B})\|_{2}\leq\epsilon$ and $\epsilon<r$; then both $\Phi_{E}(p_{B})$ and $\Phi_{E}(p_{\mathrm{adv}})$ lie in $\mathbb{B}_{r}\big(\Phi_{E}(p_{B})\big)$.
By Assumption 1 (local Lipschitz continuity of $S$),
$|S(p_{\mathrm{adv}})-S(p_{B})|\leq K\cdot\|\Phi_{E}(p_{\mathrm{adv}})-\Phi_{E}(p_{B})\|_{2}\leq K\cdot\epsilon.$
For an attack to succeed, we must have $S(p_{\mathrm{adv}})\geq\tau$, i.e.,
$S(p_{\mathrm{adv}})-S(p_{B})\geq\tau-S(p_{B}).$
Combining the two inequalities yields
$\tau-S(p_{B})\leq S(p_{\mathrm{adv}})-S(p_{B})\leq K\cdot\epsilon.$
Therefore, if $\epsilon<(\tau-S(p_{B}))/K$, it is impossible for $S(p_{\mathrm{adv}})$ to reach or exceed $\tau$. Together with the requirement $\epsilon<r$ to remain within the local Lipschitz region, we conclude that no successful adversarial example exists whenever
$\epsilon<\min\Big\{\,r,\;\frac{\tau-S(p_{B})}{K}\Big\}.$
Corollary 1 (Local Certified Radius).
The local certified robustness radius $R$ for a benign prompt $p_{B}$ in the final embedding space is
$R(p_{B})=\min\Big\{\,r,\;\frac{\tau-S(p_{B})}{K}\Big\}.$
Within the $\ell_{2}$-ball of radius $R(p_{B})$ centered at $\Phi_{E}(p_{B})$, no adversarial examples exist.
Remark. Since the Lipschitz property is local, the certification is valid only within the region where the constant $K$ upper-bounds the sensitivity. If a global bound is available (i.e., $r=\infty$), the radius simplifies to $(\tau-S(p_{B}))/K$. We use the convention that NSFW is declared whenever $S(p)\geq\tau$. This convention is consistent with the proof; alternative tie-breaking rules should be stated explicitly if used.
A.8 Details of Experimental Settings
A.8.1 Datasets
Clean Dataset. Our experiments use the validation captions of MSCOCO (Lin et al., 2014) as the clean dataset. MSCOCO is a cross-modal image-text dataset, a popular benchmark for training and evaluating T2I generation models. We remove all captions containing sensitive words to ensure the samples are benign. A total of 25,008 captions are retained.
Regular NSFW Dataset. We gather data from multiple sources to comprehensively represent various types of NSFW semantics.
  • I2P (Schramowski et al., 2023) contains 4,703 NSFW prompts sourced from real users on Lexica (Lexica, 2025). The categories include hate, harassment, violence, self-harm, sex, shocking content, and illegal activities. Specifically, we categorize the prompts into hard and soft based on their level of harmfulness.
  • 4chan Prompts (Qu et al., 2023) contain 500 NSFW prompts collected from 4chan (4chan, 2025). They predominantly encompass discriminatory and derogatory statements.
  • NSFW200 (Yang et al., 2024c) involves 200 NSFW prompts related to sexual and bloody content.
  • NSFW-LAION. We sample 1143 NSFW prompts from LAION-COCO (Schuhmann et al., 2022) to enrich datasets. These prompts mainly focus on sexual content.
A.8.2 Baselines
We select eight detection methods as baselines, encompassing both prompt-based and embedding-based approaches. These include four commercial models, two open-source models, and two SOTA methods. The commercial models are OpenAI Moderation (OpenAI, 2023; Markov et al., 2023), Azure AI Content Safety (Microsoft Azure, 2024), AWS Comprehend (Amazon Web Services, 2025), and Aliyun Text Moderation (Alibaba, 2025). These systems primarily rely on large transformer-based architectures to identify potential toxic content in prompts. The open-source moderators, NSFW-text-classifier (michellejieli, 2022) and Detoxify (unitaryai, 2022), leverage lightweight models to detect NSFW text content, offering faster inference speeds. Latent Guard (Liu et al., 2024) and GuardT2I (Yang et al., 2024b) are embedding-based methods that represent the current SOTA in detecting NSFW prompts for T2I models.
A.8.3 Target Models
We conducted experiments using three text encoders: CLIP-L, CLIP-G, and T5. Collectively, their architectures are representative of those found in the vast majority of current text-to-image models, as illustrated in Figure 6.
T2I Model Text Encoders
SD v1.4 CLIP-L
SD v1.5 CLIP-L
SD v2.1 CLIP-H
SD v3 CLIP-L, CLIP-G, T5
SD xl (Podell et al., 2023) CLIP-L, CLIP-G
FLUX.1 (Flux, 2024) CLIP-L, CLIP-G, T5
Table 6: The text encoders in different versions of open-source T2I models. The CLIP models manipulated by the attacker in the main experiment are highlighted in bold.
A.9 Multi-Categories Classifier
@@
HiddenGuard 0.9659 0.9763 0.9763 0.9821 0.9690 0.9633 0.9340

### v3

@@ IV-E The Impact of Training Data Size
ASR-H(%) ASR-M(%) CLIP Score ASR-H(%) ASR-M(%) CLIP Score
Bare 46 48 0.9597 75 79 0.9315
ESD 5 7 0.9613 10 13 0.9315
AEIOU 0 0 / 28 31 0.8850
AEIOU+ESD 0 0 / 0 0 /
@@
IV-F The Ablation Study
In our approach, we utilize NSFW features from all layers and all attention heads for detection. In this section, we will discuss the impact of using NSFW features from only a single layer of text encoder. We conduct experiments on three text encoders. Figure 7 presents the accuracy, TPR, and FPR when using each layer for detection.
In all three text encoders, even when using attention heads from a single layer, many layers still achieve high accuracy. For CLIP-L and CLIP-G, the middle layers tend to have higher accuracy, while the early and final layers show lower accuracy. Conversely, in the T5 model, the later layers exhibit higher accuracy. This highlights the distinct characteristics of the two types of text encoders.
Although using attention heads from a single layer can achieve high accuracy, we recommend using the original AEIOU method for the highest precision in detection.
V Adaptive Attack
In this section, we evaluate the robustness of AEIOU against adaptive attacks. To target our model, we design adaptive attacks based on SneakyPrompt [yang2024sneakyprompt] and MMA [yang2024mma]. They are applicable to all T2I models and can effectively bypass both internal and external safeguards. Considering their different applicable scenarios, we employ SneakyPrompt for black-box adaptive attack and MMA for white-box adaptive attack.
We utilize Stable Diffusion v1.4 [sd1.4] as the generative model, as it is the model most vulnerable to attacks. We evaluate the attack performance on the bare model, the ESD model [gandikota2023erasing], and the model employing AEIOU. The bare model only detects whether sensitive tokens are present in the prompt, while ESD fine-tunes the model to make it difficult to generate NSFW images. We assess the attacks using two metrics: attack success rate (ASR) and CLIP Score [hessel2021clipscore]. The attack success rate evaluates whether adversarial attacks can successfully generate NSFW images. The CLIP Score assesses the semantic similarity between adversarial and target prompts. The lower the CLIP score, the further the adversarial prompt deviates from the semantic meaning of the target prompt.
To ensure result reliability, both model-based and manual evaluations are conducted for attack success rate. A SOTA model [qu2023unsafe] is used for automated classification, while human evaluations are performed independently by three individuals, with the majority opinion determining the final assessment.
V-A Black-Box Adaptive Attack
In the black-box scenario, we assume the attacker has no knowledge of the model’s details but can choose prompts and query the model to obtain output. We integrate AEIOU into the text encoder of Stable Diffusion. When a potential NSFW prompt is detected, the model will refuse to generate the image. We select 100 prompts with clear NSFW semantics from NSFW200 as target prompts and use SneakyPrompt to attack the model. For all prompts successfully generated by SneakyPrompt, we then input them into the diffusion model to test whether they can generate NSFW images. Each image classified as NSFW is counted as a single success. Experimental results are shown in Table VI. ASR-H and ASR-M represent the ASR evaluated by human and classification model, respectively.
When attacking the bare model, SneakyPrompt achieves a high attack success rate. ESD significantly reduces the attack success rate, yet it cannot entirely prevent the generation of NSFW images. However, after incorporating AEIOU, SneakyPrompt is entirely thwarted and unable to generate any NSFW images. This is because SneakyPrompt only replaces a few tokens in the prompt, which does not effectively neutralize its overall meaning.
V-B White-Box Adaptive Attack
In a white-box scenario, we assume attackers can only generate images through queries. However, they possess a local copy of the text encoder identical to the target model and are aware of the AEIOU defense strategy. The attacker can target AEIOU by modifying the loss function to conduct a specific attack. The MMA attack’s loss aims to make the conditional embeddings of the adversarial prompt and the target prompt as similar as possible. To effectively attack AEIOU, we incorporate $Score(p)$ as $L_{\text{AEIOU}}$ into the original loss function. This transforms the objective of the loss function to minimize the NSFW Score while ensuring that the semantics of the adversarial prompt closely align with the target prompt.
$L=L_{\text{MMA}}+\lambda\times L_{\text{AEIOU}}$ (16)
Where $\lambda$ is a weighting factor that balances between the two components. Based on this foundation, we implement a target truncation strategy. Specifically, once $L_{\text{AEIOU}}$ exceeds the threshold by a small margin, we stop optimizing it and shift our primary focus to optimizing $L_{\text{MMA}}$. This enables adversarial prompts to approximate the semantics of the target prompt as closely as possible while avoiding detection. Consequently, the final loss is formulated as:
$L=L_{\text{MMA}}+\lambda\times\max(L_{\text{AEIOU}},\tau-\epsilon),$ (17)
Where $\tau$ is the threshold and $\epsilon$ is the margin. When updating the best prompt, we ensure that the $L_{\text{AEIOU}}$ surpasses the threshold. We conduct the attack using the default settings of MMA and evaluate it on 100 target prompts. The experimental results are presented in Table VI.
As a white-box attack, MMA demonstrates stronger capabilities than SneakyPrompt on the bare model. However, when it targets AEIOU, AEIOU exhibits robust performance, significantly reducing the attack success rate. Meanwhile, the average CLIP score of successfully generated adversarial prompts also decreases, indicating that MMA’s optimization of these prompts is not as successful as in the bare model. Moreover, although AEIOU itself does not increase the memory usage of the text encoder, the adaptive attack against it must optimize hidden states across all attention heads, significantly increasing the memory requirements. While standard MMA operates on less than 10GB of memory, adaptive MMA demands nearly 50GB, restricting its execution to commercial-grade GPUs. This makes adaptive attacks on AEIOU more challenging.
We can also combine AEIOU with other defense methods. For instance, by integrating ESD, we can reduce the success rate of MMA attacks to zero. By implementing a comprehensive defense strategy that addresses other parts of the T2I model, we can enhance the overall defensive performance and make adaptive attacks more challenging.
To further enhance AEIOU’s resilience against adaptive attacks, we incorporate adversarial prompts that successfully breach AEIOU’s defenses into the training set for additional training. The following section will provide a detailed discussion of this process.
V-C Red-Teaming Data Augmentation
We conduct red team testing using a white-box adaptive attack and collect adversarial prompts that successfully bypass AEIOU and generate NSFW images. These prompts are added to AEIOU’s training dataset for data augmentation, aiming to enhance AEIOU’s ability to resist corresponding adaptive attacks. When training, we assign them greater weight to ensure they significantly influence the model even if the sample size is limited.